Bi-prism is kept 10 cm away from the slit illuminated by monochromatic light of wavelength 5896Å. the width of the fringes obtained on screen placed at a distance of 90 cm from the biprism is( 9*10^-4)m. What is the distance between two coherent sources.
1
Expert's answer
2020-06-03T12:04:59-0400

As per the question,

distance "(a_1)= 10cm"

wavelength "(\lambda)=5896Å."

The width of the fringe obtained at the distance "(a_2)=90cm"

width of the fringe "(\beta)=9\times10^{-4}m"

distance between two co-herent source =?


"D=a_1+a_2= 10cm+90cm =100cm"


"\Rightarrow (\beta)=\frac{\lambda D}{2d}"


"\Rightarrow 2d=\frac{\lambda D}{\beta}"

"\Rightarrow 2d =\frac{5896\times 10^{-10}\times 100 \times 10^{-2}}{9\times 10^{-4}}m"

"=0.655 mm"